OpenStudy (anonymous):

Find the value of x in the matrix equation. [ 3x + 2y] = [ 6 ] -6x - 3y -6

OpenStudy (anonymous):

we multiply the top equation by two. we then add them 6x +4y = 12 -6x -3y = -6 = y=6 subsitute y into first equation 6x+24=12 6x=-12 x=-2